OpenStudy (anonymous):

The circumference of circle II is 4 feet longer than the circumference of circle I. How many feet longer is the radius of circle II than the radius of circle I?

OpenStudy (bahrom7893):

C = 2pi*r

OpenStudy (anonymous):

circumference 4 is 2pir+4

OpenStudy (bahrom7893):

C+4 = 2pi*r+4

OpenStudy (anonymous):

the radius is 2pir+4/2pi

OpenStudy (anonymous):

so the radius is r+4/2pi

OpenStudy (anonymous):

2pir

OpenStudy (anonymous):

r/2pi

OpenStudy (anonymous):

let us say radius equal 1

OpenStudy (bahrom7893):

2pir = 2pir1 + 4

OpenStudy (bahrom7893):

Solve for r1

OpenStudy (anonymous):

Circumference of Circle 2 = C2= 2Pi* R2 C2= 4+ C1 Circumference of circle 1= 2Pi*R1 therfore 4+2piR1=2pi*R2 4=2pi*(r2-r1) 2/pi +r1=r2

OpenStudy (anonymous):

1/4 pi?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

let us do this again

OpenStudy (anonymous):

so

OpenStudy (anonymous):

Circumference 2 equal 2pir+4

OpenStudy (anonymous):

radius is 2pir+4/2pi

OpenStudy (anonymous):

so radius is r+4

OpenStudy (anonymous):

your answer can also be expressed as r2-r1=0.69 (approx)

OpenStudy (anonymous):

2pir/2pi=r

OpenStudy (anonymous):

I mean

OpenStudy (anonymous):

r+2

OpenStudy (anonymous):

r+1

OpenStudy (anonymous):

ahhh

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

so the circles have different radii?

OpenStudy (anonymous):

thats the best assumption as these are not Concentric circles

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

so

OpenStudy (anonymous):

circle II is 2pir +4

OpenStudy (anonymous):

which is

OpenStudy (anonymous):

2(pir+2)?

OpenStudy (anonymous):

pir+2

OpenStudy (anonymous):

divided by pi

OpenStudy (anonymous):

r+2/pi?

OpenStudy (anonymous):

circle 2 's circumference = 2pi R2 Circle 1's circumference = 2Pi R1

OpenStudy (anonymous):

2pir

OpenStudy (anonymous):

pir

OpenStudy (anonymous):

r/pi

OpenStudy (anonymous):

why do use the same radius R, is it explicitly mentioned ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

then its better to use different variables for radii of each circle

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so

OpenStudy (anonymous):

does that help, ? then look at the solution I gave

OpenStudy (anonymous):

C=2pir1+4

OpenStudy (anonymous):

yes thats correct

OpenStudy (anonymous):

so

OpenStudy (anonymous):

C-4/2pi=r1?

OpenStudy (anonymous):

C=2pir2

OpenStudy (anonymous):

C/2pi=r2?

OpenStudy (anonymous):

lets see circumference equals 4pi

OpenStudy (anonymous):

so lets say circle 2 has radii = r2 circle one has radius = r1 circumference of circle 2 is = C2= 4+C1 BUT C2= 2*PI*r2 sub that in the equation 2*pi*r2=4+2*pi*r1 2*pi*(r2-r1)=4 (r2-r1)= 4/(2*pi) r2-r1=2/pi r2=r1+2/pi

OpenStudy (anonymous):

does that help??

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

I'll take another practice test and see you guys later